7. Roundtree Software is going public using an auction IPO. The firm has received the following bids: Price (S) 4.00 13.80 13.60 3.40 13.20 13.00 2.80 Number of Shares 100,000 200,000 500,000 000,000 ,200,000 800,000 400,000
Assuming Roundtree would like to sell 1.8 million shares in its IPO, what will be the winning auction offer price?
Answer #1

The price is set in such a way that an investment bank should be able to sell all the available shares and at the highest price available in the market.
Number of available shares =1.8*1000000=1800000

When the price is set at $13.4, total number of shares that will be sold=100000+200000+500000+1000000=1800000. This is exactly the number of available shares.

If the price is set above $13.4 at $13.6, the number of shares that will be sold=100000+200000+500000=800000. At this price, all the available shares (that is 1.8 million) will not be sold. So, this price will be rejected.

As all the shares can be sold at a price of $13.4, any price less than $13.4 will be rejected.

Winning auction offer price will be equal to $13.4